if the measure of one of two complementary angles is 22 more than three times the measure of the other, find the measure of the angle
Answer by stanbon(75887)

if the measure of one of two complementary angles is 22 more than three times the measure of the other, find the measure of the angle
-------
1st angle x
2nd angle 3x+22
--------------------
Equation:
x + 3x+22 = 90
4x = 68
x = 17 degrees
---------------------
3x+22 = 3*17+22 = 73 degrees
---------------------------------
Cheers,
Stan H.
=============